A man distributed Php 1,500,000 among his three children to start a business of their own. The eldest received PHp 100,000 less than the sum of the shares of the other two, while the second child received Php 200,000 more than the share of the youngest child. How much did each receive?
Answer by TimothyLamb(4379)

x = amount to oldest child
y = amount to middle child
z = amount to youngest child
---
x + y + z = 1500000
x = y + z - 100000
y = z + 200000
---
put the system of linear equations into standard form
---
x + y + z = 1500000
x - y - z = -100000
y - z = 200000
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = amount to oldest child = 700,000
y = amount to middle child = 500,000
z = amount to youngest child = 300,000
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---